OpenStudy (anonymous):

9+2/3x=81 (please show work)

OpenStudy (anonymous):

2/3x = 81-9 (subtract 9 from both sides) 2/3x = 72 (simplify) x= 108 (multiply 3/2 from both sides) (you can also multiply both sides by 3 and then divide by 2. either way will give you the correct answer)